Feed aggregator

COMMENT ON B91 and B95 Radiological Waste Container Relocation

Recent Log Entries -

Log Number: 3858250Logbook(s): Body: 

Update from Adam Hartbmerger when asked about progress by MRI, the mechanical subcontractor:

They were able to get the hole drilled through the wall in Bldg 91 today. The current plan is to drill at Bldg 95 on Monday. Hopeful to work on the piping on Tuesday, pending any developments on the LCW pumps pulling them away from 91/95.

COMMENT ON Injector Upgrade - Phase 1 Installation (Gun to Chopper)

Recent Log Entries -

Log Number: 3858232Logbook(s): Body: 

Week 9 Wrap-Up (Day #45 couched within):
AIPINJ --
* After alignment, four new pedestals grouted in
* Ceiling lifting plate for AIPINJ Phase II Booster install
opportunistically installed and load tested

* Y-chamber --
** Viewers' screen support plates beveled, cleaned
** Viewer assemblies installed in Y-chamber
** Viewer screens mounted and aligned
** NEG pump tower installed

* Wien --
** V-wien magnet field mapping completed and results
reviewed
** H-wien magnet field mapping begun
** V-wien electrodes installed in chem-cleaned vacuum
chamber and gap surveyed for

iocsofthahfm restored

Recent Log Entries -

Log Number: 3858230Logbook(s): Tags: Body: restorelib 2.2.1 Target restore date was 2020-10-03 11:46:18 The output from the myRestore execution was as follows: iocsofthahfm system@2020-10-03 11:46 2(2) restored ha_hfm@2020-10-03 11:46 2(2) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ced -A -s -i iocsofthahfm (success) Procedure return status: Success

HDIce run3 delayed, plan going forward

Recent Log Entries -

Log Number: 3858210Logbook(s): Body: 

Carlos and Marcy will attach a turbo pump to the gun and heat the line overnight. Then after HDIce is done with NMR studies tomorrow, Carlos will krypton process. Once field emission cleans up, we want to soak under vacuum conditions at 200 kV. Then we remove dummy puck8 and insert puck31 with bulk GaAs, and hopefully we don't wake up another field emitter in the process. Beam by Monday? more later

FE from gun, won't sit at 200 kV anymore

Recent Log Entries -

Log Number: 3858196Logbook(s): Entry Makers: poelkerBody: 

Gun tripping OFF, lots of field emission, lots of FE on viewer. It seems likely this is a result of difficulty inserting the puck into the cathode electrode.

Time to Kr process, then heat and reactivate the photocathode, sigh.

iocse4 restored

Recent Log Entries -

Log Number: 3858193Logbook(s): Tags: Body: restorelib 2.2.1 Target restore date was 2020-11-20 11:30:28 The output from the myRestore execution was as follows: iocse4 system@2020-11-20 11:30 2(2) restored seebpm@2020-11-20 11:30 2641(2641) restored SEE_BURT 61(61) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ced -A -s -i iocse4 (success) Procedure return status: Success

Pages

Subscribe to cebaf2013 aggregator